In a electricle sell comparison of eletromotive froce is 1.5voltand internai resistance is 2 ohm.with in this sell 1 ohm,2 ohm,10 ohm resistance is connect in series combination.s what will be potential different and lost volt?
1
Expert's answer
2020-04-13T10:05:57-0400

As all the resistances are in series hence the total resistance will be (2+1+2+10) = 15 ohm

Hence current in the circuit will be ="\frac V R=\frac{1.5}{15}=0.1 A"

Hence the voltage drop across the the internal resistance of 2 ohm =0.1 x 2=0.2 volt